On driving from destination A to B, the following probabilities were calculated. The probability to reach destination B on time is 0.8, the probability to reach destination B 1 hour late is 0.5, the probability to reach destination B 2 hours late is 0.1, and the probability to reach destination B 3 hours late is 0.08. What is the expected value of lateness?
Answer #1

Given,

Probability to reach destination on time = P(x=0) = 0.8

Probability to reach destination 1 hour late = P(x=1) = 0.5

Probability to reach destination 2 hours late = P(x=2) = 0.1

Probability to reach destination 3 hours late = P(x=3) = 0.08

xi 0 1 2 3
P(xi) 0.8 0.5 0.1 0.08

Expected value of lateness E(x) is computed as follows: